National Defence Academy (NDA) · Topic Collection

Quantitative Aptitude PYQs

Arithmetic, number systems, percentages, ratios, profit and loss, time and work, algebra, geometry, mensuration, and trigonometry.

2520 Total Questions21 Official PapersYears: 2015–2025Search in Quantitative Aptitude
Showing questions 1901–1920 of 2520Page 96 of 100
NDA Mathematics 10 Sept 2017 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1901

National Defence Academy (NDA) 2017• Mathematics• Mathematics

A committee of two persons is selected from two men and two women. The probability that the committee will have exactly one woman is

  1. AOption (a)

    \(\frac{1}{6}\)

  2. BOption (b)

    \(\frac{2}{3}\)

  3. COption (c)

    \(\frac{1}{3}\)

  4. DOption (d)

    \(\frac{1}{2}\)

View answer & explanation

Correct answer

Answer: (b)B.

\(\frac{2}{3}\)

Why this answer is correct

Concept:

Probability: It is a measure of likelihood of an event to occur.

Probability of an event to occur, P(E) \(= \frac{{{\rm{favourable\;outcomes}}}}{{{\rm{total\;outcomes}}}}\)

Note:

Combination is the number of ways in which we can select a group of objects from a set. For example, selecting r things from set of n things is given as nCr.

 

Calculation:

Number of ways for selecting one man and one woman = 2C1 × 2C1

Number of ways for selecting two persons from group of four = 4C2

Probability of committee having one man and one woman \(= \frac{{{{\rm{\;}}^2}{{\rm{C}}_1}{\rm{\;}} \times {{\rm{\;}}^2}{{\rm{C}}_1}}}{{{{\rm{\;}}^4}{{\rm{C}}_2}}}\)

\(= \frac{{2 \times 2}}{{\frac{{4 \times 3}}{2}}}\)

= 2/3

NDA Mathematics 10 Sept 2017 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1902

National Defence Academy (NDA) 2017• Mathematics• Mathematics

Let a die be loaded in such a way that even faces are twice likely to occur as the odd faces. What is the probability that a prime number will show up when the die is tossed?

  1. AOption (a)

    \(\frac{1}{3}\)

  2. BOption (b)

    \(\frac{2}{3}\)

  3. COption (c)

    \(\frac{4}{9}\)

  4. DOption (d)

    \(\frac{5}{9}\)

View answer & explanation

Correct answer

Answer: (c)C.

\(\frac{4}{9}\)

Why this answer is correct

Properties of Probability:

  • Probability of an event, 0 ≤ P(E) ≤ 1
  • The Sum of the probability of all outcomes is equal to one. Let A and B be two events then:

 

P(A) + P(B) = 1

Calculation:

Sample space = {1, 2, 3, 4, 5, 6}

Let the probability of face coming up to be odd be p.

So, the probability of face coming up to be even be 2p.

So, we know sum of probability of all outcomes is equal to one.

P(odd face) + P(even face) = 1

⇒ P(1) + P(3) + P(5) + P(2) + P(4) + P(6)

⇒ x + x + x + 2x + 2x + 2x = 1

⇒ 9x = 1

⇒ x = 1/9

Now, the prime numbers in sample space are 2, 3, 5

So, probability that a prime number will show up = P(2) + P(3) + P(5)

= 2x + x + x

= 4x

= 4/9

NDA Mathematics 10 Sept 2017 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1903

National Defence Academy (NDA) 2017• Mathematics• Mathematics

Let the sample space consist of non-negative integers up to 50, X denote the numbers which are multiples of 3 and Y denote the odd numbers. Which of the following is/are correct?

  1. \({\rm{P}}\left( {\rm{X}} \right) = \frac{8}{{25}}\) 2. \({\rm{P}}\left( {\rm{Y}} \right) = \frac{1}{2}\)

Select the correct answer using the code given below.

  1. AOption (a)

    1 only

  2. BOption (b)

    2 only

  3. COption (c)

    Both 1 and 2

  4. DOption (d)

    Neither 1 nor 2

View answer & explanation

Correct answer

Answer: (d)D.

Neither 1 nor 2

Why this answer is correct

Concept:

Probability: It is a measure of likelihood of an event to occur.

Probability of an event to occur, P(E) \(= \frac{{{\rm{number\;of\;favourable\;outcomes}}}}{{{\rm{number\;of\;total\;outcomes}}}}\)

 

Calculation:

Let,

S: Sample space consists of non-negative number up to 50.

X: Numbers which are multiple of 3

Y: Numbers which are odd.

S = {0, 1, 2, …, 50} & n(S) = 51

X = {3, 6, 9, …, 48} & n(X) = 16

Y = {1, 3, 5, …, 49} & n(Y) = 25

Probability of an event, P(E) \(= \frac{{{\rm{number\;of\;favourable\;outcomes}}}}{{{\rm{number\;of\;total\;outcomes}}}} = \frac{{{\rm{n}}\left( {\rm{E}} \right)}}{{{\rm{n}}\left( {\rm{S}} \right)}}\)

P(X) = 16/51

P(Y) = 25/51

Hence, **both statement

(1) & (2)** are incorrect.

NDA Mathematics 10 Sept 2017 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1904

National Defence Academy (NDA) 2017• Mathematics• Mathematics

For two events A and B, let \({\rm{P}}\left( {\rm{A}} \right) = \frac{1}{2},{\rm{\;P}}\left( {{\rm{A}} \cup {\rm{B}}} \right) = \frac{2}{3}\) and \({\rm{P}}\left( {{\rm{A}} \cap {\rm{B}}} \right) = \frac{1}{6}\). What is the P(A̅ ∩ B) equal to?

  1. AOption (a)

    \(\frac{1}{6}\)

  2. BOption (b)

    \(\frac{1}{4}\)

  3. COption (c)

    \(\frac{1}{3}\)

  4. DOption (d)

    \(\frac{1}{2}\)

View answer & explanation

Correct answer

Answer: (a)A.

\(\frac{1}{6}\)

Why this answer is correct

Concept:

Formulas used:

  • P(A ∪ B) = P(A) + P(B) - P(A ∩ B)
  • P(A̅ ∩ B) = P(B) - P(A ∩ B)

 

Calculation:

Given: \({\rm{P}}\left( {\rm{A}} \right) = \frac{1}{2},{\rm{\;P}}\left( {{\rm{A}} \cup {\rm{B}}} \right) = \frac{2}{3}\) and \({\rm{P}}\left( {{\rm{A}} \cap {\rm{B}}} \right) = \frac{1}{6}\)

To find: P(A̅ ∩ B)

We know that, P(A ∪ B) = P(A) + P(B) - P(A ∩ B)

So, P(B) = P(A ∪ B) + P(A ∩ B) - P(A)

\(\Rightarrow {\rm{P}}\left( {\rm{B}} \right) = \frac{2}{3} + \frac{1}{6} - \frac{1}{2}\)

⇒ P(B) = 2/6

⇒ P(B) = 1/3

 

P(A̅ ∩ B) = P(B) - P(A ∩ B)

⇒ P(A̅ ∩ B) = 1/3 - 1/6

⇒ P(A̅ ∩ B) = 1/6

NDA Mathematics 10 Sept 2017 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1905

National Defence Academy (NDA) 2017• Mathematics• Mathematics

Consider the following

Statements:

  1. Coefficient of variation depends on the unit of measurement of the variable.
  2. Range is measure of dispersion.
  3. Mean deviation is least when measured about median.

Which of the above statements are correct?

  1. AOption (a)

    1 and 2 only

  2. BOption (b)

    2 and 3 only

  3. COption (c)

    1 and 3 only

  4. DOption (d)

    1, 2 and 3

View answer & explanation

Correct answer

Answer: (b)B.

2 and 3 only

Why this answer is correct

Concept:

Coefficient of variation: It is defined as ratio of standard deviation to mean.

\({\rm{CV}} = \frac{{{{\rm{\sigma }}_{\rm{X}}}}}{{{\rm{\bar X}}}}\)

Range: It is difference between largest value and smallest value.

Mean deviation: Let x1, x2, …, xn be n observations, then:

Mean deviation \(= \frac{{\mathop \sum \nolimits_{{\rm{i}} = 1}^{\rm{n}} \left| {{{\rm{x}}_{\rm{i}}} - {\rm{\bar x}}} \right|}}{{\rm{n}}}\)

It is minimum when measure around mean.

 

Calculation:

The coefficient of variation is sometimes referred to the standard

deviation because the value of the coefficient of variation is

independent of the unit of measurement scale

So, statement (1) is incorrect.

Range is difference between largest value and smallest value. So, it defines how much the value is spread or dispread.

So, statement (2) is correct.

Mean deviation is minimum when measured around median.

So, statement (3) is correct.

NDA Mathematics 10 Sept 2017 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1906

National Defence Academy (NDA) 2017• Mathematics• Mathematics

Given that the arithmetic mean and standard deviation of a sample of 15 observations are 24 and 0 respectively. Then which one of the following is the arithmetic mean of the smallest five observations in the data?

  1. AOption (a)

    0

  2. BOption (b)

    8

  3. COption (c)

    16

  4. DOption (d)

    24

View answer & explanation

Correct answer

Answer: (d)D.

24

Why this answer is correct

Concept:

Standard deviation: Let x1, x2, …, xn be n observations of X, then:

\({\rm{\sigma }} = \sqrt {\frac{{\mathop \sum \nolimits_{{\rm{i}} = 1}^{\rm{n}} {{\left( {{{\rm{x}}_{\rm{i}}} - {\rm{\bar x}}} \right)}^2}}}{{\rm{n}}}} \)

Calculation:

Given:

Number of observations (n) = 15, Mean (x̅) = 24 & Standard deviation (σ) = 0

So, \({\rm{\sigma }} = \sqrt {\frac{{\mathop \sum \nolimits_{{\rm{i}} = 1}^{\rm{n}} {{\left( {{{\rm{x}}_{\rm{i}}} - {\rm{\bar x}}} \right)}^2}}}{{\rm{n}}}} \)

\(\Rightarrow 0 = \sqrt {\frac{{\mathop \sum \nolimits_{{\rm{i}} = 1}^{\rm{n}} {{\left( {{{\rm{x}}_{\rm{i}}} - {\rm{\bar x}}} \right)}^2}}}{{\rm{n}}}} \)

Squaring both sides, we get,

\(\Rightarrow 0 = \frac{{\mathop \sum \nolimits_{{\rm{i}} = 1}^{\rm{n}} {{\left( {{{\rm{x}}_{\rm{i}}} - {\rm{\bar x}}} \right)}^2}}}{{\rm{n}}}\)

\(\Rightarrow \mathop \sum \limits_{{\rm{i}} = 1}^{\rm{n}} {\left( {{{\rm{x}}_{\rm{i}}} - {\rm{\bar x}}} \right)^2} = 0\)

⇒ (x1 - x̅)2 + (x2 - x̅)2 + … + (x15 - x̅)2 = 0

So, every deviation is squared and summed and summation is zero. This means every deviation is individually zero.

⇒ (xi - x̅)2 = 0 for i = 1, 2, …, 15

⇒ xi = x̅ for i = 1, 2, …, 15

⇒ xi = 24 for i = 1, 2, …, 15

 So, mean of smallest five observations \(= \frac{{24 + 24 + 24 + 24 + 24}}{5}\)

= 24

NDA Mathematics 10 Sept 2017 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1907

National Defence Academy (NDA) 2017• Mathematics• Mathematics

Which one of the following can be considered as appropriate pair of values of regression coefficient of y on x and regression coefficient of x on y?

  1. AOption (a)

    (1, 1)

  2. BOption (b)

    (-1, 1)

  3. COption (c)

    \(\left( { - \frac{1}{2},\;2} \right)\)

  4. DOption (d)

    \(\left( {\frac{1}{3},\frac{{10}}{3}} \right)\)

View answer & explanation

Correct answer

Answer: (a)A.

(1, 1)

Why this answer is correct

Concept:

  • The product of regression coefficient of y on x and regression coefficient of x on y is always less than or equal to 1. So \({{\rm{b}}_{{\rm{xy}}}} \times {{\rm{b}}_{{\rm{yx}}}} \le 1\) where bxy and byx are coefficient of regression.
  • The signs of both coefficients should be same.

 

Calculation:

We know that the signs of both coefficients should be same. 

So option 2 and 3 can’t be answer because both coefficient are of different sign.

We will check 1 option and 4 option because both have same sign so we will check product of both coefficient which should be less than equal to 1.

Checking 4th option

\( \Rightarrow \frac{1}{3} \times \frac{{10}}{3} \le 1\)

\( \Rightarrow \frac{{10}}{9} \le 1\)

\(\Rightarrow 1.11 \le 1\) Not true so it can’t be answer

Checking 1st option

⇒ 1 × 1 ≤ 1

\( \Rightarrow 1 \le 1\) It is true so it is the right answer

NDA Mathematics 10 Sept 2017 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1908

National Defence Academy (NDA) 2017• Mathematics• Mathematics

Let A and B be two events with \({\rm{P}}\left( {\rm{A}} \right) = \frac{1}{3},{\rm{\;P}}\left( {\rm{B}} \right) = \frac{1}{6}\) and \({\rm{P}}\left( {{\rm{A}} \cap {\rm{B}}} \right) = \frac{1}{{12}}\). What is the P(B|A̅) equal to?

  1. AOption (a)

    \(\frac{1}{5}\)

  2. BOption (b)

    \(\frac{1}{7}\)

  3. COption (c)

    \(\frac{1}{8}\)

  4. DOption (d)

    \(\frac{1}{10}\)

View answer & explanation

Correct answer

Answer: (c)C.

\(\frac{1}{8}\)

Why this answer is correct

Concept:

Conditional probability: It is the measure of probability of an event A is occurring given that another event B is already occurred, then:

\({\rm{P}}({\rm{B}}|{\rm{A}}) = \frac{{{\rm{P}}\left( {{\rm{A}} \cap {\rm{B}}} \right)}}{{{\rm{P}}\left( {\rm{A}} \right)}}\)

Formulas used:

  • P(A̅ ∩ B) = P(B) - P(A ∩ B)
  • P(A̅ ) = 1 - P(A)

 

Calculation:

Given: \({\rm{P}}\left( {\rm{A}} \right) = \frac{1}{3},{\rm{\;P}}\left( {\rm{B}} \right) = \frac{1}{6}\) and \({\rm{P}}\left( {{\rm{A}} \cap {\rm{B}}} \right) = \frac{1}{{12}}\)

To find: P(B|A̅)

We know that, conditional probability, \({\rm{P}}\left( {{\rm{A|B}}} \right) = \frac{{{\rm{P}}\left( {{\rm{A}} \cap {\rm{B}}} \right)}}{{{\rm{P}}\left( {\rm{B}} \right)}}\)

So, \({\rm{P}}\left( {{\rm{B|\bar A}}} \right) = \frac{{{\rm{P}}\left( {{\rm{\bar A}} \cap {\rm{B}}} \right)}}{{{\rm{P}}\left( {{\rm{\bar A}}} \right)}}\)

So, P(A̅ ∩ B) = P(B) - P(A ∩ B)

⇒ P(A̅ ∩ B) = 1/6 - 1/12

⇒ P(A̅ ∩ B) = 1/12

Now, P(A̅ ) = 1 - P(A)

⇒ P(A̅) = 1 - 1/3

⇒ P(A̅) = 2/3

∴ \({\rm{P}}\left( {{\rm{B|\bar A}}} \right) = \frac{{\frac{1}{{12}}}}{{\frac{2}{3}}}{\rm{\;}}\)

⇒ P(B|A̅) = 1/8

NDA Mathematics 10 Sept 2017 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1909

National Defence Academy (NDA) 2017• Mathematics• Mathematics

In a binomial distribution, the mean is \(\frac{2}{3}\) and the variance is \(\frac{5}{9}\). What is the probability that X = 2?

  1. AOption (a)

    \(\frac{5}{36}\)

  2. BOption (b)

    \(\frac{25}{36}\)

  3. COption (c)

    \(\frac{25}{216}\)

  4. DOption (d)

    \(\frac{25}{54}\)

View answer & explanation

Correct answer

Answer: (c)C.

\(\frac{25}{216}\)

Why this answer is correct

Concept:

Binomial distribution: If a random variable X has binomial distribution as B (n, p) with n and p as parameters, then the probability of random variable is given as:

P( X = k) = nCk pk (1 - p)(n - k)

Where, n is number of observations, p is the probability of success & (1 - p) is probability of failure.

 

Properties:

  • Mean of the distribution (μX) = n × p
  • The variance (σ2x) = n × p × (1 - p)
  • Standard deviation (σx) = √{np(1 - p)}

 

Calculation:

Given:

Mean = n × p = 2/3    ….

(1) Variance = n × p × (1 - p) = 5/9    …. (2) Dividing equation (1) by equation (2)

\(\frac{{{\rm{np}}\left( {1 - {\rm{p}}} \right)}}{{{\rm{np}}}} = \frac{{\frac{5}{9}}}{{\frac{2}{3}}}\)

⇒ (1 - p) = 5/6

⇒ p = 1/6

Put p = 1/6 in equation (1)

n × (1/6) = 2/3

⇒ n = 4

Now, P(X = 2) = 4C2 × (1/6)2 × (1 - 1/6)(4 - 2)

P(X = 2) = 6 × (1/6)2 × (5/6)2

\(\Rightarrow {\rm{P}}\left( {{\rm{x}} = 2} \right) = \frac{{25}}{{{6^4}}}\)

\(\Rightarrow {\rm{P}}\left( {{\rm{X}} = 2} \right) = \frac{{25}}{{216}}\)

NDA Mathematics 10 Sept 2017 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1910

National Defence Academy (NDA) 2017• Mathematics• Mathematics

The probability that a ship safely reaches a part is \(\frac{1}{3}\). The probability that out of 5 ships, at least 4 ships would arrive safely is

  1. AOption (a)

    \(\frac{1}{243}\)

  2. BOption (b)

    \(\frac{10}{243}\)

  3. COption (c)

    \(\frac{11}{243}\)

  4. DOption (d)

    \(\frac{13}{243}\)

View answer & explanation

Correct answer

Answer: (c)C.

\(\frac{11}{243}\)

Why this answer is correct

Concept:

Binomial distribution: If a random variable X has binomial distribution as B (n, p) with n and p as parameters, then the probability of random variable is given as:

P(X = k) = nCk pk (1 - p)(n - k)

Where, n is number of observations, p is the probability of success & (1 - p) is probability of failure.

 

Calculation:

Given:

Success of a ship to reach the port safely = p = 1/3

Probability that out of 5 ships, at least 4 should arrive = P(X = 4) + P(X = 5)

= 5C4 (1/3)4 (1 - 1/3)(5 - 4) + 5C5 (1/3)5 (1 - 1/3)(5 - 5)

\(= 5 \times \frac{1}{{81}} \times \frac{2}{3} + 1 \times \frac{1}{{243}} \times 1\)

\(= \frac{{10}}{{243}} + \frac{1}{{243}}\)

= 11/243

NDA Mathematics 10 Sept 2017 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1911

National Defence Academy (NDA) 2017• Mathematics• Mathematics

What is the probability that at least two persons out of a group of three persons were born in the same month (disregard year)?

  1. AOption (a)

    \(\frac{{33}}{{144}}\)

  2. BOption (b)

    \(\frac{{17}}{{72}}\)

  3. COption (c)

    \(\frac{{1}}{{144}}\)

  4. DOption (d)

    \(\frac{{2}}{{9}}\)

View answer & explanation

Correct answer

Answer: (b)B.

\(\frac{{17}}{{72}}\)

Why this answer is correct

Calculation:

Let us assume that each month is equally likely as a birth month.

A: Probability that all three were born in same month.

B: Probability two of them were born in same month.

E: Probability that at least two born in same month

P(E) = P(A) + P(B)

 

So, there are 123 choices for three birthdays all together, takin in order

For all three born in same month, there are 12 choices.

\({\rm{P}}\left( {\rm{A}} \right) = \frac{{12}}{{12 \times 12 \times 12}}\)

For 2 born in same month and the one in another; selecting any two born in any 12 months and the left one in another month.

\({\rm{P}}\left( {\rm{B}} \right) = \frac{{{{\rm{\;}}^3}{{\rm{C}}_2} \times 12 \times 11}}{{12 \times 12 \times 12}}\)

\({\rm{P}}\left( {\rm{E}} \right) = \frac{{12}}{{12 \times 12 \times 12}} + \frac{{{{\rm{\;}}^3}{{\rm{C}}_2} \times 12 \times 11}}{{12 \times 12 \times 12}}\)

\(= \frac{1}{{144}} + \frac{{33}}{{144}}\)

= 34/144

= 17/72

NDA Mathematics 10 Sept 2017 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1912

National Defence Academy (NDA) 2017• Mathematics• Mathematics

It is given that X̅ = 10, Y̅ = 90, σX = 3, σY = 12 and rXY = 0.8. The regression equation of X on Y is

  1. AOption (a)

    Y = 3.2X + 58

  2. BOption (b)

    X = 3.2Y + 58

  3. COption (c)

    X = - 8 + 0.2Y

  4. DOption (d)

    Y = - 8 + 0.2X

View answer & explanation

Correct answer

Answer: (c)C.

X = - 8 + 0.2Y

Why this answer is correct

Concept:

Regression equation: Let x and y be two variables, then equation is given as:

\(\frac{{{\rm{x}} - {\rm{\bar x}}}}{{{{\rm{\sigma }}_{\rm{x}}}}} = {\rm{r}}\left[ {\frac{{{\rm{y}} - {\rm{\bar y}}}}{{{{\rm{\sigma }}_{\rm{y}}}}}} \right]\)

 

Calculation:

Given: x̅ = 10, y̅ = 90, σx = 3, σy = 12 and rxy = 0.8

So, regression equation is given as,

\(\frac{{{\rm{x}} - {\rm{\bar x}}}}{{{{\rm{\sigma }}_{\rm{x}}}}} = {\rm{r}}\left[ {\frac{{{\rm{y}} - {\rm{\bar y}}}}{{{{\rm{\sigma }}_{\rm{y}}}}}} \right]\)

\(\Rightarrow \frac{{{\rm{x}} - 10}}{3} = 0.8\left[ {\frac{{{\rm{y}} - 90}}{{12}}} \right]\)

⇒ x - 10 = 0.2 × {y - 90}

⇒ x = 0.2y - 18 + 10

⇒ x = 0.2y - 8

⇒ X = 0.2Y - 8

NDA Mathematics 10 Sept 2017 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1913

National Defence Academy (NDA) 2017• Mathematics• Mathematics

If \(P\left( B \right) = \frac{3}{4},\;P\left( {A \cap B \cap \overline{C}} \right) = \frac{1}{3}\) and \(P\;\left( {\overline{A}\; \cap \;B\; \cap \;\overline{C}} \right) = \frac{1}{3},\) then what is P (B ∩ C) equal to?

  1. AOption (a)

    \(\frac{1}{{12}}\)

  2. BOption (b)

    \(\frac{3}{{4}}\)

  3. COption (c)

    \(\frac{1}{{15}}\)

  4. DOption (d)

    \(\frac{1}{{9}}\)

View answer & explanation

Correct answer

Answer: (a)A.

\(\frac{1}{{12}}\)

Why this answer is correct

Calculation:

Given that

\(\Rightarrow \;P\left( B \right) = \frac{3}{4},\;P\left( {A \cap B \cap {C^c}} \right) = \frac{1}{3}\) And \(P\;\left( {{A^c} \cap B \cap \;{C^c}} \right) = \frac{1}{3}\)

Form the above diagram we can write

\(\Rightarrow {\rm{P}}\left( {\rm{B}} \right) = {\rm{P}}\left( {{{\rm{A}}^{\rm{c}}}{\rm{\;}} \cap {\rm{B\;}} \cap {{\rm{C}}^{\rm{c}}}} \right) + {\rm{P}}\left( {{\rm{A\;}} \cap {\rm{B\;}} \cap {{\rm{C}}^{\rm{c}}}} \right) + {\rm{P}}\left( {{\rm{A\;}} \cap {\rm{B\;}} \cap {\rm{C}}} \right) + {\rm{P}}\left( {{{\rm{A}}^{\rm{c}}}{\rm{\;}} \cap {\rm{B}} \cap {\rm{C}}} \right)\)

\(\Rightarrow {\rm{P}}\left( {{\rm{B}} \cap {\rm{C}}} \right) = {\rm{\;P}}\left( {{\rm{A\;}} \cap {\rm{B\;}} \cap {\rm{C}}} \right) + {\rm{P}}\left( {{{\rm{A}}^{\rm{c}}}{\rm{\;}} \cap {\rm{B}} \cap {\rm{C}}} \right)\)

\( \Rightarrow {\rm{P}}\left( {\rm{B}} \right) = {\rm{P}}\left( {{{\rm{A}}^{\rm{c}}}{\rm{\;}} \cap {\rm{B\;}} \cap {{\rm{C}}^{\rm{c}}}} \right) + {\rm{P}}\left( {{\rm{A\;}} \cap {\rm{B\;}} \cap {{\rm{C}}^{\rm{c}}}} \right) + {\rm{P}}\left( {{\rm{B\;}} \cap {\rm{C}}} \right)\)

\(\Rightarrow {\rm{P}}\left( {{\rm{B}} \cap {\rm{C}}} \right) = {\rm{P}}\left( {\rm{B}} \right) - {\rm{P}}\left( {{{\rm{A}}^{\rm{c}}}{\rm{\;}} \cap {\rm{B\;}} \cap {{\rm{C}}^{\rm{c}}}} \right) - {\rm{P}}\left( {{\rm{A\;}} \cap {\rm{B\;}} \cap {{\rm{C}}^{\rm{c}}}} \right)\)

Put all given values

\(\Rightarrow {\rm{P}}\left( {{\rm{B}} \cap {\rm{C}}} \right) = \frac{3}{4} - \frac{1}{3} - \frac{1}{3}\)

\(\Rightarrow {\rm{P}}\left( {{\rm{B}} \cap {\rm{C}}} \right) = \frac{3}{4} - \frac{2}{3}\)

\(\therefore {\rm{P}}\left( {{\rm{B}} \cap {\rm{C}}} \right) = \frac{1}{{12}}\)

NDA Mathematics 10 Sept 2017 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1914

National Defence Academy (NDA) 2017• Mathematics• Mathematics

The following table gives the monthly expenditure of two families:

Expenditure (in Rs.)
ItemsFamily AFamily B
Food3,5002,700
Clothing500800
Rent1,5001,000
Education2,0001,800
Miscellaneous2,5001,800

 

In constructing a pie diagram to the above data, the radii of the circles are to be chosen by which one of the following ratios?

  1. AOption (a)

    1 : 1

  2. BOption (b)

    10 : 9

  3. COption (c)

    100 : 91

  4. DOption (d)

    5 : 4

View answer & explanation

Correct answer

Answer: (b)B.

10 : 9

Why this answer is correct

Concept:

Pie Chart: A pie chart is a circular statistical graphic, which is divided into slices to illustrate numerical proportion.

Some properties:

  • The arc length and central of each slide is proportional to the quantity it represents.
  • The area of the pie chart denoted the total quantity.

 

Calculation:

Area of the circle = Total expenditure

Total expenditure of family A = Rs. 10000

Area of circle A = 10000

Total expenditure of family B = Rs. 8100

Area of circle B = 8100

\(\frac{{{\rm{Area\;of\;cirlce\;A}}}}{{{\rm{Area\;of\;circle\;B}}}} = \frac{{10000}}{{8100}}\)

\(\Rightarrow \frac{{{\rm{\pi r}}_{\rm{a}}^2}}{{{\rm{\pi r}}_{\rm{b}}^2}} = \frac{{100}}{{81}}\)

\(\Rightarrow \frac{{{{\rm{r}}_{\rm{a}}}}}{{{{\rm{r}}_{\rm{b}}}}} = \frac{{10}}{9}\)

⇒ ra : rb = 10 : 9

NDA Mathematics 10 Sept 2017 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1915

National Defence Academy (NDA) 2017• Mathematics• Mathematics

If a variable takes values 0, 1, 2, 3, …, n with frequencies 1, C(n, 1), C(n, 2), C(n, 3),…, C(n, n) respectively, then the arithmetic mean is

  1. AOption (a)

    2n

  2. BOption (b)

    n + 1

  3. COption (c)

    n

  4. DOption (d)

    \(\frac{{\rm{n}}}{2}\)

View answer & explanation

Correct answer

Answer: (d)D.

\(\frac{{\rm{n}}}{2}\)

Why this answer is correct

Concept:

Mean: Let X be a random variable with values xi for i = 1, 2, …, n with frequencies fi. So:

Mean \(= \frac{{\mathop \sum \nolimits_{{\rm{i}} = 0}^{\rm{n}} {{\rm{x}}_{\rm{i}}}{{\rm{f}}_{\rm{i}}}}}{{\mathop \sum \nolimits_{{\rm{i}} = 0}^{\rm{n}} {{\rm{f}}_{\rm{i}}}}}\)

Binomial expansion formula:

  • (1+ x)n = nC0x0 + nC1x1 + nC2x2 + … + nCnxn

 

Calculation:

Given: 

xi = i for i = 0, 1, 2, …, n

fi = nCi for i = 0, 1, 2, …, n

So, mean \(= \frac{{\mathop \sum \nolimits_{{\rm{i}} = 0}^{\rm{n}} {{\rm{x}}_{\rm{i}}}{{\rm{f}}_{\rm{i}}}}}{{\mathop \sum \nolimits_{{\rm{i}} = 0}^{\rm{n}} {{\rm{f}}_{\rm{i}}}}}\)

\(= \frac{{0 \times {{\rm{\;}}^{\rm{n}}}{{\rm{C}}_0} + 1 \times {{\rm{\;}}^{\rm{n}}}{{\rm{C}}_1} + 2 \times {{\rm{\;}}^{\rm{n}}}{{\rm{C}}_2} + \ldots + {\rm{n}} \times {{\rm{\;}}^{\rm{n}}}{{\rm{C}}_{\rm{n}}}}}{{{{\rm{\;}}^{\rm{n}}}{{\rm{C}}_0} + {{\rm{\;}}^{\rm{n}}}{{\rm{C}}_1} + {{\rm{\;}}^{\rm{n}}}{{\rm{C}}_2} + \ldots + {{\rm{\;}}^{\rm{n}}}{{\rm{C}}_{\rm{n}}}}}\)

Now we know that,

(1+ x)n = nC0x0 + nC1x1 + nC2x2 + … + nCnxn ----(1)

Differentiating both sides w.r.t x

⇒ n(1 + x)n-1 = 0 + 1 × nC1x + 2 × nC2x1 + … + n × nCnxn ----(2)

Put x = 1 in equation (2)

⇒ n × 2n-1 = 0 + 1 × nC1 + 2 × nC2 + … + n × nCn ----(3)

Put x = 1 in equation (1)

⇒ 2n = nC0x0 + nC1x1 + nC2x2 + … + nCnxn ----(4)

 

So, \(\frac{{0 \times {{\rm{\;}}^{\rm{n}}}{{\rm{C}}_0} + 1 \times {{\rm{\;}}^{\rm{n}}}{{\rm{C}}_1} + 2 \times {{\rm{\;}}^{\rm{n}}}{{\rm{C}}_2} + \ldots + {\rm{n}} \times {{\rm{\;}}^{\rm{n}}}{{\rm{C}}_{\rm{n}}}}}{{{{\rm{\;}}^{\rm{n}}}{{\rm{C}}_0} + {{\rm{\;}}^{\rm{n}}}{{\rm{C}}_1} + {{\rm{\;}}^{\rm{n}}}{{\rm{C}}_2} + \ldots + {{\rm{\;}}^{\rm{n}}}{{\rm{C}}_{\rm{n}}}}} = \frac{{{\rm{n}} \times {2^{{\rm{n}} - 1}}}}{{{2^{\rm{n}}}}}\) [Using equation (3) & (4)]

\(= \frac{{\rm{n}}}{2}\)

NDA Mathematics 10 Sept 2017 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1916

National Defence Academy (NDA) 2017• Mathematics• Mathematics

In a multiple-choice test, an examinee either knows the correct answer with probability p, or guesses with probability 1 - p. The probability of answering a question correctly is \(\rm 1\over m\) , if he or she merely guesses. If the examinee answers a question correctly, the probability that he or she really knows the answer is

  1. AOption (a)

    \(\frac{{{\rm{mp}}}}{{1 + {\rm{mp}}}}\)

  2. BOption (b)

    \(\frac{{{\rm{mp}}}}{{1 + \left( {{\rm{m}} - 1} \right){\rm{p}}}}\)

  3. COption (c)

    \(\frac{{\left( {{\rm{m}} - 1} \right)}}{{1 + \left( {{\rm{m}} - 1} \right){\rm{p}}}}\)

  4. DOption (d)

    \(\frac{{\left( {{\rm{m}} - 1} \right){\rm{p}}}}{{1 + {\rm{mp}}}}\)

View answer & explanation

Correct answer

Answer: (b)B.

\(\frac{{{\rm{mp}}}}{{1 + \left( {{\rm{m}} - 1} \right){\rm{p}}}}\)

Why this answer is correct

Concept:

Probability: It is a measure of likelihood of an event to occur.

Probability of an event to occur, P(E) \(= \frac{{{\rm{favourable\;outcomes}}}}{{{\rm{total\;outcomes}}}}\)

 

Calculation:

Probability of knowing correct answer = p

Probability of guessing the correct answer \(= \left( {1 - {\rm{p}}} \right) \times \frac{1}{{\rm{m}}}\)

So, probability of answering correctly \(= {\rm{p}} + \frac{{1 - {\rm{p}}}}{{\rm{m}}}\)

Probability of answering correctly considering that, he or she knows the correct answer \(= \frac{{\rm{p}}}{{{\rm{p}} + \frac{{1 - {\rm{p}}}}{{\rm{m}}}}}\)

\(= \frac{{{\rm{mp}}}}{{{\rm{mp}} + 1 - {\rm{p}}}}\)

\(= \frac{{{\rm{mp}}}}{{1 + \left( {{\rm{m}} - 1} \right){\rm{p}}}}\)

NDA Mathematics 10 Sept 2017 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1917

National Defence Academy (NDA) 2017• Mathematics• Mathematics

If x1 and x2 are positive quantities, then the condition for the difference between the arithmetic mean and the geometric mean to be greater than 1 is

  1. AOption (a)

    \({{\rm{x}}_1} + {{\rm{x}}_2} > 2\sqrt {{{\rm{x}}_1}{{\rm{x}}_2}} \)

  2. BOption (b)

    \(\sqrt {{{\rm{x}}_1}} + \sqrt {{{\rm{x}}_2}} > \sqrt 2 \)

  3. COption (c)

    \(\left| {\sqrt {{{\rm{x}}_1}} - \sqrt {{{\rm{x}}_2}} } \right| > \sqrt 2 \)

  4. DOption (d)

    \({{\rm{x}}_1} + {{\rm{x}}_2} < 2{\rm{\;}}\left( {\sqrt {{{\rm{x}}_1}{{\rm{x}}_2}} + 1} \right)\)

View answer & explanation

Correct answer

Answer: (c)C.

\(\left| {\sqrt {{{\rm{x}}_1}} - \sqrt {{{\rm{x}}_2}} } \right| > \sqrt 2 \)

Why this answer is correct

Concept:

Arithmetic mean: It is sum of all observations upon number of observations.

If x1, x2, x3, …., xn are the observations, then mean is given as:

\({\rm{\bar X}} = \frac{{{{\rm{x}}_1} + {{\rm{x}}_2} + \ldots + {{\rm{x}}_{\rm{n}}}}}{{\rm{n}}}\)

Geometric mean: It is the nth root of product n numbers. Let x1, x2, …, xn be n values of X, then:

Geometric mean \({{\rm{\bar X}}_{{\rm{geom}}}} = \sqrt[{\rm{n}}]{{\mathop \prod \limits_{{\rm{i}} = 1}^{\rm{n}} {{\rm{x}}_{\rm{i}}}}}\)

Calculation:

Given: x1 & x2 are positive.

A.M. - G.M. > 1

So, \(\frac{{{{\rm{x}}_1} + {{\rm{x}}_2}}}{2} - \sqrt {{{\rm{x}}_1}{{\rm{x}}_2}} > 1\)

\(\Rightarrow {\rm{\;}}\frac{{{{\rm{x}}_1} + {{\rm{x}}_2}}}{2} > \sqrt {{{\rm{x}}_1}{{\rm{x}}_2}} + 1\)

⇒ x1 + x2 > 2√(x1x2) + 2

⇒ x1 + x2 - 2√(x1x2) > 2

\(\Rightarrow {\left( {\sqrt {{{\rm{x}}_1}} - \sqrt {{{\rm{x}}_2}} } \right)^2} > 2\)

\(\Rightarrow \left| {\sqrt {{{\rm{x}}_1}} - \sqrt {{{\rm{x}}_2}} } \right| > \sqrt 2 \)

NDA Mathematics 10 Sept 2017 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1918

National Defence Academy (NDA) 2017• Mathematics• Mathematics

Consider the following

Statements:

  1. Variance is unaffected by change of origin and change of scale.
  2. Coefficient of variance is independent of the unit of observations.

Which of the statements given above is/are correct?

  1. AOption (a)

    1 only

  2. BOption (b)

    2 only

  3. COption (c)

    Both 1 and 2

  4. DOption (d)

    Neither 1 nor 2

View answer & explanation

Correct answer

Answer: (b)B.

2 only

Why this answer is correct

Concept:

  • \({{\rm{\sigma }}^2} = \mathop \sum \limits_{i = 1}^N \frac{{{{\left( {{x_i} - \;\mu } \right)}^2}}}{N}\)
  • \(\mu = \sum \frac{{{x_i}}}{N}\)

Where, σ2 is variance, μ is Mean and N is Total numbers of observation

  • The coefficient of variation expresses the standard deviation as a percentage of the sample mean. \(\Rightarrow {\bf{CV}} = \frac{{\bf{\sigma }}}{{\bf{\mu }}} \times 100\)

Where, σ is standard deviation and μ is mean of population

Calculation:

  1. Variance is unaffected by change of origin and change of scale.
  • Variance is independent of change of origin as the change in origin is uniformly added to all the values and hence the mean also and hence, when \({\left( {{{\rm{x}}_{\rm{i}}} - {\rm{\mu }}} \right)^2}\) is calculated, it remains unaffected.

For example: x1 + 5, x2 + 5 ……… xn+5 than mean and xi both will change in same manner so while we calculate \({\left( {{{\rm{x}}_{\rm{i}}} - {\rm{\mu }}} \right)^2}\) it won’t change.

  • But, change of scale alters all values unevenly and hence, variance changes.

For example: 2x1, 2x2 …………. 2xn than mean and xi both will change so variance also change.

So that first statement is not completely right

  1. Coefficient of variance is independent of the unit of observations.
  • The coefficient of variation expresses the standard deviation as a percentage of the sample mean. \(CV = \frac{\sigma }{\mu } \times 100\)

Both σ, μ have same unit so it gets cancel out.

So, CV is independent of the unit of variance 2nd statement is right.

NDA Mathematics 10 Sept 2017 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1919

National Defence Academy (NDA) 2017• Mathematics• Mathematics

Five sticks of length 1, 3, 5, 7 and 9 feet are given. Three of these sticks are selection at random. What is the probability that the selected sticks can form a triangle?

  1. AOption (a)

    0.5

  2. BOption (b)

    0.4

  3. COption (c)

    0.3

  4. DOption (d)

    0

View answer & explanation

Correct answer

Answer: (c)C.

0.3

Why this answer is correct

Concept:

Condition of triangle: Let a, b, c be the length of sides of the triangle, then:

Sum of smaller sides should be greater than third side, i.e., a + b > c

 

Probability: It is a measure of likelihood of an event to occur.

Probability of an event to occur, P(E) \(= \frac{{{\rm{favourable\;outcomes}}}}{{{\rm{total\;outcomes}}}}\)

Note:

Combination is the number of ways in which we can select a group of objects from a set. For example, selecting r things from set of n things is given as nCr

 

Calculation:

Given: five length of sticks 1, 3, 5, 7, 9 feet

Condition of a triangle to be formed, if a, b, c be length of sides of triangle then:

a + b > c

Total number of cases = 5C3 = 10

Cases from which triangles can be formed:

  1. 3 + 5 > 7
  2. 3 + 7 > 9
  3. 5 + 7 > 9

 

Number of favourable outcomes = 3

So, probability that triangle will be formed = 3/10 = 0.3

NDA Mathematics 10 Sept 2017 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1920

National Defence Academy (NDA) 2017• Mathematics• Mathematics

The coefficient of correlation when coefficients of regression are 0.2 and 1.8 is

  1. AOption (a)

    0.36

  2. BOption (b)

    0.2

  3. COption (c)

    0.6

  4. DOption (d)

    0.9

View answer & explanation

Correct answer

Answer: (c)C.

0.6

Why this answer is correct

Concept:

Coefficient of Correlation: The geometric mean between two regression coefficient is equal to coefficient of correlation i.e.,

\({\rm{r}} = \sqrt {{{\rm{b}}_{{\rm{yx}}}}{{\rm{b}}_{{\rm{xy}}}}} \)

Some other properties of regression coefficient:

  • The regression coefficient of y on x is represented by byx and x on y as bxy.
  • Both regression coefficient must have same sign.
  • If one regression coefficient is greater than unity than others will be lesser than unity.
  • Arithmetic mean of both regression coefficient:
  • \(\frac{{{{\rm{b}}_{{\rm{yx}}}} + {{\rm{b}}_{{\rm{xy}}}}}}{2} \ge {\rm{r}}\)

 

Calculation:

Given: byx = 0.2 & bxy = 1.8

To find: coefficient of correlation (r)

So, coefficient of correlation = r

\(= \sqrt {0.2 \times 1.8} \)

= √(0.36)

= 0.6

How should I prepare for Quantitative Aptitude in National Defence Academy (NDA)?

Solving previous year questions by topic helps identify recurring concepts, question patterns, and frequently asked sub-topics. Attempt each question, review the step-by-step explanation, and note down key formulas or facts for quick revision before the exam.